OpenStudy (anonymous):

fine b and c so that y=7x^2+bx+c has a vertex (4,-10)

OpenStudy (campbell_st):

well the line of symmetry can be found using \[x = \frac{-b}{2a}\] from the vertex you know the like of symmetry is x = 4 and from the equation you know a = 7 so you have the equation \[4 = \frac{-b}{2 \times 7}\] solve for b. now substitute b into the original equation... to find c you know the vertex lies on the parabola. So substitute x = 4 and y = -10 into the equation and solve for c. hope it helps